# Robotic deployment on construction sites: considerations for safety and productivity impact

### v1

@@ I INTRODUCTION
The imperative of ensuring human safety in shared workplaces with robotic systems presents a major challenge for the field deployment of such systems. The risks associated with human-robot collaboration encompass both direct and indirect hazards, with direct risks arising from collisions between the platform and obstacles such as equipment or personnel, while indirect risks stem from environmental factors affecting robot functionality, including noise and wind, as well as potential distractions posed by the presence of the robot itself. Addressing these concerns, Moud, Hashem Izadi, et al. proposed a model for evaluating the safety of mobile robot deployment based on physical characteristics and environmental considerations [1].
In a similar vein, Augustsson et al. advocated for the establishment of defined safety zones surrounding robots to facilitate safe interaction with humans, with the flexibility of zone geometries tailored to accommodate varying work scenarios [2]. Complementing this approach, Truong et al. emphasized the importance of human safety in shared environments, implementing a system wherein robots dynamically generate safe navigation zones upon detecting individuals through sensor feedback [3]. Furthermore, Shin et al. proposed a framework aligned with ISO/TS 15066 to evaluate the collision peak pressure of collaborative robots in proximity to humans, underlining the criticality of safety considerations within shared workspaces [4].
Our work focuses on the construction industry, recognized for its inherent hazards to human workers. It presents a particularly challenging environment for the deployment of autonomous mobile robots. Kim et al. highlighted persistent security concerns surrounding human-robot collisions in construction sites, accentuated by the mobility of robots [5]. While early endeavors in human-robot collaboration within construction settings have demonstrated potential efficiency gains, the complexities of maintaining safety alongside productivity remain significant [6].
The objective of this paper is to study the safety and productivity impacts of the deployment of autonomous mobile robots on construction sites, as perceived in a real industry context. For this purpose, we first review the industry standards on this subject, followed by directed interviews with construction industry experts in mobile robotics. Finally, we propose a mobile robot prototype specifically equipped for construction site progress monitoring.
@@ II Industry Standards
Among the most recent standards, ANSI/RIA R15.08 stands as a pivotal framework for regulating industrial mobile robots (IMRs). Within its purview, both autonomous mobile robots (AMRs) and Automated Guided Vehicles (AGVs) are encompassed under the IMR designation when deployed within industrial settings. However, AGVs and AMRs exhibit distinct operational scopes: while AGVs adhere to predetermined routes, AMRs leverage sensor technologies to navigate around obstacles autonomously. Rooted in foundational principles from ANSI/RIA R15.06 and ANSI/ITSDF B56.5, which respectively address industrial robot safety and guided industrial vehicles, ANSI/RIA R15.08 stipulates comprehensive requirements for IMRs and their associated attachments. Noteworthy provisions include specifications for operational modes, control functionalities, and presence-sensing devices, among others.
In tandem with adherence to industry standards, due consideration must also be accorded to local regulations pertinent to the targeted workplace environment. National and local codes and regulations supersede industry standards, mandating strict compliance, particularly in construction sites where Health and Safety Standards hold paramount importance. Each construction site is overseen by a Safety Agent tasked with ensuring adherence to these regulations.
On a construction site, the traffic plan and traffic control are essential, as they constitute a measure to prevent unexpected contact between pedestrians and vehicles. To this end, the project manager must plan the movement of the latter to restrict backing maneuvers and put in place safety measures to protect any person circulating on the site. They must also inform, beforehand, any person who has to circulate on the site of the safety measures provided for in the circulation plan (CSTC, art. 2.8 and 2.8.2). This traffic plan must obviously include any IMR likely to be deployed. The site signaller must meet certain conditions, in particular: wear high-visibility fluorescent yellow-green safety clothing, remain visible to vehicle drivers, and stay out of the latter’s path (CSTC, art. 2.8.4). Therefore, any worker must remain visible and identifiable to an IMR by wearing high visibility clothing.
Self-propelled vehicles must be equipped with effective brakes and warning devices to be used when approaching pedestrians, doors, turns, and dangerous locations. This requirement does not apply to mechanical rams mounted on crawlers, skidders, and all-terrain vehicles (CSTC, art. 3.10.2). All equipment must only be used by an experienced driver or under their supervision (CSTC, art. 3.10.4). Finally, all self-propelled vehicles must be equipped with an audible warning device which must be installed within reach of the driver, have a sound specific to the type of equipment and different from any other signal from the same construction site and have sufficient noise level to dominate construction noise (CSTC, art. 3.10.12).
@@ III Insights from the industry
Thanks to our close collaboration with two major companies in Canada already deploying robotic systems on their construction sites, we tackle some of the main challenges through directed interviews.
@@ IV-B The Fields
  • Second Building, Early Construction Stage: This early-stage construction site was littered with construction materials and equipment. We deployed the robot on the second floor, where worker presence was minimal, in a vast yet cluttered navigation space. Piles of construction material and assorted obstacles demanded agile maneuvering of the robot.
  • Third Building, Middle stage: This factory plant was in the middle stages of construction, with partially installed production lines and scattered construction materials. The larger spaces offered ample room for the robot to navigate easily.
@@ V CONCLUSIONS
Ensuring human safety in shared workplaces with robotic systems presents This paper presented our endeavors to study safety considerations and productivity impacts of deploying autonomous mobile robots in construction sites. Thanks to our collaborations with major construction companies in Canada we obtained insights on the subject from the industry perspective and performed deployments in real scenarios, allowing us to identify key points and strategies that we believe are very helpful in these situations. The main key points are: performing a thorough assessment of the site before the deployment, to identify possible challenges for the robot; testing the robotic system in simulations and physical environments that recreate the main characteristics and challenges of the real construction site; maintaining checklists to be verified after each deployment to avoid mistakes and increase consistency. In the future, these protocols can be applied with other robotic platforms and extended for more varied environments.

### v2

@@ I INTRODUCTION
The imperative of ensuring human safety in workplaces shared with robotic systems presents a significant challenge for the field deployment of such systems. Risks associated with human-robot collaboration encompass both direct and indirect hazards. Direct risks arise from collisions between the platform and obstacles, such as equipment or personnel, while indirect risks stem from environmental factors that affect robot functionality, including noise and wind, as well as potential distractions posed by the presence of the robot itself. Addressing these concerns, Moud, Hashem Izadi, et al. proposed a model for evaluating the safety of mobile robot deployment based on physical characteristics and environmental considerations [1].
Similarly, Augustsson et al. advocated for the establishment of defined safety zones around robots to facilitate safe interactions with humans, with the flexibility of zone geometries tailored to accommodate varying work scenarios [2]. Complementing this approach, Truong et al. emphasized the importance of human safety in shared environments by implementing a system wherein robots dynamically generate safe navigation zones upon detecting individuals through sensor feedback [3]. Additionally, Shin et al. proposed a framework aligned with ISO/TS 15066 to evaluate the peak pressure of collisions involving collaborative robots in proximity to humans, underlining the critical importance of safety considerations within shared workspaces [4].
Our work focuses on the construction industry, recognized for its inherent hazards to human workers. It presents a particularly challenging environment for the deployment of autonomous mobile robots. Kim et al. highlighted persistent safety concerns surrounding human-robot collisions at construction sites, accentuated by the mobility of the robots [5]. While early endeavors in human-robot collaboration within construction settings have demonstrated potential efficiency gains, the complexities of maintaining safety alongside productivity remain significant [6].
The objective of this paper is to study the safety and productivity impacts of deploying autonomous mobile robots on construction sites, as perceived in a real industry context. To this end, we first review the industry standards on this subject, followed by directed interviews with construction industry experts in mobile robotics. Finally, we share our experience at deploying a mobile robot prototype specifically equipped for monitoring progress at construction sites.
@@ II Industry Standards
Among the most recent standards, ANSI/RIA R15.08 serves as a pivotal framework for regulating industrial mobile robots (IMRs). This standard encompasses both autonomous mobile robots (AMRs) and automated guided vehicles (AGVs) within industrial settings. While AGVs follow predetermined routes, AMRs leverage sensor technologies to navigate autonomously around obstacles. Rooted in principles from ANSI/RIA R15.06, which addresses industrial robot safety, and ANSI/ITSDF B56.5, which pertains to guided industrial vehicles, ANSI/RIA R15.08 outlines comprehensive requirements for IMRs, including operational modes, control functionalities, and presence-sensing devices.
In addition to industry standards, compliance with local regulations, such as the Safety Code for the construction industry [^1] https://www.legisquebec.gouv.qc.ca/en/document/cr/s-2.1,%20r.%204 in Quebec, Canada, is critical, particularly in construction environments where health and safety standards are paramount. Local and national codes often supersede industry standards, mandating strict adherence. On construction sites, a Safety Agent oversees compliance with these regulations to ensure worker safety.
Traffic control is essential on construction sites to prevent collisions between pedestrians and vehicles. The project manager must develop a traffic plan to restrict backing maneuvers and implement safety measures to protect individuals on the site. This plan should include provisions for any IMRs deployed. Site signalers must wear high-visibility clothing, remain visible to vehicle drivers, and stay out of their path. Consequently, workers must also be visible and identifiable to IMRs by wearing high-visibility clothing.
Self-propelled vehicles must be equipped with effective brakes and warning devices, used when approaching pedestrians, doors, turns, and hazardous locations. These requirements ensure that vehicles are operated safely, especially in the dynamic environment of a construction site. Only experienced drivers, or those under supervision, should operate such equipment. Additionally, all self-propelled vehicles must have an audible warning device within the driver’s reach, distinct from other site signals, and loud enough to be heard over construction noise.
These regulatory measures not only ensure the operational efficiency of mobile robots but also facilitate their safe integration into the dynamic and complex environments of construction sites. However, adhering to standards alone is insufficient to guarantee the safety of these robotic deployments; it is equally important to secure the acceptance of the technology by the workers.
@@ III Insights from the industry
Thanks to our close collaboration with major companies in Canada already deploying robotic systems on their construction sites, we tackle some of the main challenges through directed interviews.
@@ IV-B The Fields
  • Commercial Building, Early Construction Stage: This early-stage construction site was littered with construction materials and equipment. We deployed the robot on the second floor, where worker presence was minimal, in a vast yet cluttered navigation space. Piles of construction material and assorted obstacles demanded agile maneuvering of the robot.
  • Industrial Building, Middle stage: This factory plant was in the middle stages of construction, with partially installed production lines and scattered construction materials. The larger spaces offered ample room for the robot to navigate easily.
@@ V CONCLUSIONS
Ensuring human safety in shared workplaces with robotic systems presents significant challenges. This paper has explored the safety considerations and productivity impacts of deploying autonomous mobile robots on construction sites. Through our collaborations with leading construction companies in Canada, we have gained valuable industry insights and conducted deployments in real-world scenarios. This experience has enabled us to identify particularly effective strategies. These strategies include performing a thorough site assessment prior to deployment to anticipate potential challenges, testing the robotic system in simulations and environments that closely mimic the actual construction site (BIM-powered), and maintaining checklists to ensure consistency and prevent errors after each deployment. Additionally, maintaining a strong connection with workers to ensure their acceptance of the technology and focusing on functional aesthetics rather than anthropomorphic features have proven essential. Looking ahead, these protocols could be adapted for use with other robotic platforms and applied in a broader range of environments.